OpenStudy (anonymous):

Use the three steps to solve the problem. Two planes leave an airport at noon. If the eastbound plane flies at 560 mph and the westbound plane flies at 500 mph, at what time will they be 2,000 miles apart? Round your answer to the nearest minute.

OpenStudy (anonymous):

560x + 500x = 2000 1060x = 2000 x = 2000/1060 = 1.88 hrs 0.88 hours = 52.8 minutes Answer: 1 hour and 53 minutes

OpenStudy (anonymous):

Hope it helps!